### v1

# Do Not (Always) Look Right: Investigating the Capabilities of Decoder-Based Large Language Models for Sequence Labeling

Abstract
Pre-trained language models based on masked language modeling (MLM) objective excel in natural language understanding (NLU) tasks. While fine-tuned MLM-based encoders consistently outperform causal language modeling decoders of comparable size, a recent trend of scaling decoder models to multiple billion parameters resulted in large language models (LLMs), making them competitive with MLM-based encoders. Although scale amplifies their prowess in NLU tasks, LLMs fall short of SOTA results in information extraction (IE) tasks, many framed as sequence labeling (SL). However, whether this is an intrinsic limitation of LLMs or whether their SL performance can be improved remains unclear. To address this, we explore strategies to enhance the SL performance of “open” LLMs (Llama2 and Mistral) on IE tasks. We investigate bidirectional information flow within groups of decoder blocks, applying layer-wise removal or enforcement of the causal mask (CM) during LLM fine-tuning. This approach yields performance gains competitive with SOTA SL models, matching or outperforming the results of CM removal from all blocks. Our findings hold for diverse SL tasks, proving that “open” LLMs with layer-dependent CM removal outperform strong MLM-based encoders and instruction-tuned LLMs. However, we observe no effect from CM removal on a small scale when maintaining an equivalent model size, pre-training steps, and pre-training and fine-tuning data.
Do Not (Always) Look Right: Investigating the Capabilities of Decoder-Based Large Language Models for Sequence Labeling
David Dukić  Jan Šnajder TakeLab, Faculty of Electrical Engineering and Computing, University of Zagreb {david.dukic, jan.snajder}@fer.hr
1 Introduction
Pre-trained language models (PLMs) built upon the transformer architecture have demonstrated exceptional performance across many natural language understanding (NLU) tasks (Radford et al., 2018; Devlin et al., 2019; Yang et al., 2019; Clark et al., 2020; Raffel et al., 2020). Typically, achieving SOTA results in tasks such as sequence classification and sequence labeling involved a two-step process: pre-training on unlabeled corpora, followed by fine-tuning on task-specific data – a process often referred to as transfer learning (Raffel et al., 2020). Two prevailing architectures emerged, each coupled with a compatible pre-training paradigm: (1) the decoder architecture, utilizing causal language modeling (CLM) as its pre-training objective, and (2) the encoder architecture, with masked language modeling (MLM) pre-training objective.
In direct comparisons with a similar number of model parameters and employing a transfer learning strategy, MLM-based encoders consistently outperformed CLM-based decoders on NLU tasks (Devlin et al., 2019). However, a shift occurred within the NLP community when encoder models ceased being scaled to the same magnitude of parameters and pre-training data as their decoder counterparts. Instead, there has been a pronounced trend toward scaling decoder models to multiple billion parameters, which led to large language models (LLMs). LLMs boosted the performance of many NLU tasks by combining their text generation capabilities with prompting techniques, eliminating the need for fine-tuning model parameters (Liu et al., 2023).
Nevertheless, there is still considerable room for improvement, even for the largest decoder models, such as ChatGPT, which falls far behind SOTA results on fundamental NLU tasks. This holds particularly for information extraction (IE) tasks (Han et al., 2023), such as named entity recognition (NER), aspect-based sentiment analysis (ABSA), and event extraction (EE). These tasks are often cast as sequence labeling (token classification) tasks, and tackling sequence labeling (SL) by prompting LLMs has been proven quite difficult (Wang et al., 2023a). However, it is unclear whether the subpar performance of LLMs on SL tasks is their innate limitation or if it can be remedied. As the community stopped scaling up encoders, LLMs are solidifying their position as the de facto standard in the field. Thus, improving the SL performance of LLMs stands as the sole viable option.
While an investigation of LLM capabilities for SL has practical merit, the challenge is that many LLMs are locked behind paywalls, and their weights are not publicly accessible. Fortunately, the community has been training and publicly releasing LLMs with multiple billions of parameters, such as Llama2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and OPT (Zhang et al., 2022). Leveraging parameter-efficient fine-tuning techniques and quantization, including QLoRA (Dettmers et al., 2023), facilitates experimentation with fine-tuning these models on task-specific data.These strategies allow leveraging “open” LLM’s hidden states for task-specific classifiers, enabling fine-tuning on consumer-grade hardware with comparable or better performance than full fine-tuning.
Using the hidden state of the decoder’s last token serves as a reliable feature for classifying the input sequence. However, for SL, the attempts to utilize each token’s hidden state as a feature do not yield satisfying results. The causal mask, integral to endowing decoders with text generation capabilities, limits bidirectional information flow, preventing the model from “looking right,” i.e., attending to tokens in the sequence positioned to the right of the current token. For many SL tasks, the token’s label depends on the succeeding tokens, and omitting this context often results in subpar performance. Recently, Li et al. (2023) found that completely removing the causal mask from the Llama2 model during fine-tuning substantially enhances SL performance on NER.
In this paper, we explore strategies to enhance the SL performance of “open” LLMs on IE tasks. Building on insight from Li et al. (2023), we explore bidirectional information flow within different decoder blocks. We find that layer-wise causal mask removal across LLM blocks can result in gains competitive with SOTA models, contrasting the approaches of removing or keeping the causal mask in all layers. Our analysis on a series of IE tasks (NER, ABSA, and EE) shows that layer-wise “looking right” is highly task-dependent. We compare against strong encoder-based sequence taggers and instruction-tuned LLMs. “Open” LLMs without causal masks in specific layers outperform these baselines in almost all scenarios. Finally, we draw comparisons with encoder-based models, assessing the relative strengths of the decoder and encoder architectures regarding parameter scale, training data, and the particular SL task. To rigorously evaluate the architectures, we pre-train small encoder and decoder models from scratch with an identical number of parameters on the same data. Our findings indicate the superiority of the same scale encoder architecture over the decoder on SL tasks, which is consistent even when the causal mask is removed while fine-tuning the small pre-trained decoder models. This suggests that the observed SL performance boost in LLMs upon removing the causal mask stems from scaling up.
2 Background and Related Work
Encoders and Decoders.
The first decoder-based PLM with CLM pre-training that gained traction was GPT (Radford et al., 2018), while the first encoder-based PLM, with grand MLM pre-training success, was BERT (Devlin et al., 2019). Pre-training with CLM teaches the model to generate text and does not allow it to use the sequence context right to the token being processed. This is enabled via a causal mask. While beneficial and necessary for coherent text generation, restricting the model to look right is detrimental if we want to use the decoder to produce token-level embeddings. Encoders pre-trained with MLM and without the causal mask are much more successful in solving SL tasks because of the enabled bidirectional information flow. Although decoder-based models are bad at labeling tokens in a sequence, they can be used as successful sequence classifiers where the representation of the last token in the sequence is standardly leveraged for classification. Many decoder models implement sequence classification heads on top of decoder blocks, but not SL heads (Wolf et al., 2020). Li et al. (2023) found that removing the causal mask from all decoder layers of Llama2 (Touvron et al., 2023) improves the performance across sequence classification tasks and the IE task of named entity recognition (NER). However, Li et al. (2023) do not examine in depth why this phenomenon occurs, whether it occurs with layer-wise causal mask removal, and whether it is limited to the NER task. We build on this and inspect the effect of different unmasking configurations across layers of models with Llama decoder blocks, yielding interesting findings for IE and SL tasks. Furthermore, we train smaller decoders with CLM and encoders with MLM to demonstrate that the superiority of unmasked Llama2 is a consequence of the model and training data scale. The authors of Megatron-LM (Shoeybi et al., 2019) showed that scaling the BERT model to four billion parameters with MLM and the GPT-2 model to eight billion parameters with CLM brings gains on NLU tasks. Nevertheless, the community stopped scaling encoders and continued scaling decoders. This shift in strategy has given rise to notable properties in decoder-based LLMs, including enhanced zero- and few-shot learning capabilities (Wei et al., 2021; Brown et al., 2020), the ability for in-context learning (ICL) (Brown et al., 2020; Liu et al., 2023), and chain-of-thought prompting (Wei et al., 2022). Examples of LLMs with these emerging properties are GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2023), and ChatGPT.
Universal IE.
Many IE tasks, such as NER, subtasks of ABSA, and EE, can effectively be treated as SL tasks. Traditionally, these tasks were addressed with encoder models, mapping each token’s hidden state to task label space (Zhang et al., 2020; Zhou and Chen, 2021; Pontiki et al., 2015). However, this approach necessitates training a separate model for each task. More recent methods aim to infer the underlying structure of multiple IE tasks simultaneously and develop universal information extraction (UIE) systems, employing encoder, decoder, or encoder-decoder architectures. Often, the models of choice for this approach are encoder-decoder PLMs such as BART (Lewis et al., 2020) and T5 (Raffel et al., 2020). Building on T5 and instead of training task-specific classifiers, Paolini et al. (2021) propose the TANL framework, treating IE tasks as translation tasks between augmented natural languages. Similarly, Lu et al. (2022) introduce UIE, a text-to-structure generation framework leveraging T5. The framework generates target extractions through a schema-based prompt mechanism, with the structured schema learned in an unsupervised fashion. Further, the authors of LasUIE (Fei et al., 2022) leverage syntactic structure information for UIE improvement. Meanwhile, Wang et al. (2022a) extend the GLM (Du et al., 2022) with structure pre-training, proposing DeepStruct and reformulating structure prediction as a series of triple prediction tasks. Recently, Ping et al. (2023) revealed UniEX, a paradigm for UIE, utilizing BERT-like encoders to convert text-based IE into a token-pair problem. This is achieved by modifying the attention mechanism to extract spans in a unified manner. Another encoder-based approach recasts IE tasks into a multi-span cyclic graph extraction problem to end up with a non-autoregressive graph decoding algorithm, proving beneficial for zero and few-shot settings (Zhu et al., 2023). Finally, Lou et al. (2023) reconceptualize IE as a triplet prediction problem through semantic matching for UIE. We argue that the benefits of having a unified IE framework are undeniable. However, including more IE tasks or the same tasks with domain-specific data into each framework would still require re-training the UIE extractors. Therefore, we stick to the one model per IE and SL task approach, relying on parameter-efficient methods that yield compact and extensible models (Houlsby et al., 2019).
In-context Learning for IE.
IE tasks can be tackled with ICL, relying on the emergent properties of LLMs. Nevertheless, Pang et al. (2023) report that the performance of ICL lags behind SOTA supervised expert models. They develop a model that learns to guide the prompt so that LLMs output better results. Another example is GPT-NER (Wang et al., 2023a), where the authors transform the NER task into a generation one utilizing ICL and ChatGPT.
Instruction Tuning for IE.
Releasing the weights of encoder-decoder and decoder-only pre-trained LLMs paved the way for instruction tuning (IT) (Mishra et al., 2022). This approach involves fine-tuning the model to address specific tasks in a supervised generative manner, providing the model with instructions for each data point. The authors of InstructionNER (Wang et al., 2022b) propose a framework for low-resource NER. They reformulate the task as a generation problem with better prompt templates and instruction-tune the T5 model, stating that current prompt templates are primarily designed for sentence-level tasks and inappropriate for SL objectives. Extending the fusion of UIE and IT, InstructUIE (Wang et al., 2023b) reveals that multi-task IT can give results comparable to BERT (Devlin et al., 2019) in a supervised setting. Finally, Scaria et al. (2023) find that instruction-tuned encoder-decoder PLMs excel at ABSA subtasks. They used the Tk-instruct model, which is the PLM T5 fine-tuned in a supervised manner on various instruction tasks (Wang et al., 2022c).
3 Layer Group Unmasking
The causal mask is crucial in decoders, preventing the model from attending to future tokens and facilitating autoregressive text generation. This constraint is implemented by adding the causal triangular mask, denoted as CM, to the dot product of the query and key attention matrices. The resulting sum is passed through the softmax function in the scaled dot-product attention mechanism, as introduced by Vaswani et al. (2017). Formally, CM and the scaled dot-product attention are given as follows:
[…]
Here, $Q$, $K$, and $V$ represent the query, key, and value attention matrices, respectively, and $d_{k}$ is the dimension of queries and keys. Effectively, applying softmax for the tokens with value $\infty$- ∞ in the CM results in attention scores of 00. To allow the LLM to “sometimes look right” (hence the title of our paper), we strategically choose in which decoder blocks to replace all $\infty$- ∞ entries in the CM with 00. This equates to the CM removal (actually, we leave the positions of padding tokens to $\infty$- ∞, but replace all others with 00). As Li et al. (2023) observed, removing CM across all decoder layers of Llama2 during fine-tuning increases the model’s performance on the NER SL task by a large margin. This is surprising because the model was pre-trained with the restriction to “look right” and can learn to attend to future tokens by backpropagating on the training data over a few epochs when the CM is removed. Motivated by this finding, we try to boost the performance of “open” LLMs further, trying out combinations of bidirectional information flow enablement (“unmasking”) or restriction (“masking”) per layer groups consisting of $b$ decoder blocks. We refer to these decisions as layer group “unmasking configurations.” Considering that Llama2 is made up of $n=32$ decoder blocks, and each block can either remove or enforce CM, we deal with a search space of $2^{n}=2^{32}$ possibilities. To limit the search space, we group the $n$ decoder blocks into $m=4$ layer groups with $b=8$ decoder blocks per layer group. Here, we only consider unmasking or masking all blocks per layer group. This leaves us with $162^{m}=$ 16 configurations per LLM and task. We encode our unmasking configurations with binary four-digit codes ranging from $0000$ (all masked) to $1111$ (all unmasked) where each 00 in the code represents a layer group with $b$ masked blocks, and vice-versa, each $1$ represents a layer group with $b$ unmasked blocks. These unmasking configurations are interpreted left-to-right, where the first digit in the code refers to the first $b$ decoder blocks closest to the model’s input.
4 Experimental Setup
4.1 Datasets
We focus our experiments on IE tasks that can be framed as SL. Specifically, we choose three IE tasks: NER, ABSA with aspect term extraction and polarity subtasks, and a subtask of EE called trigger classification. Furthermore, since we tackle SL tasks, we include one more standard SL task unrelated to IE – text chunking. Dataset statistics are given in Table 1.
Dataset Train Validation Test
#Sent #Sent #Sent
CoNLL03 14041 3250 3453
ACE05 14672 873 711
ABSA Restaurants 2737 304 800
Table 1: Statistics for the datasets and their splits.
NER and Chunking.
For NER and text chunking, we choose CoNLL03, a standard and widely used benchmark (Tjong Kim Sang and De Meulder, 2003). We use a dataset version from “Hugging Face Datasets” (Lhoest et al., 2021) with an IOB2 sequence tagging scheme, which has train, validation, and test splits.
Aspect Term Extraction and Polarity.
We use the dataset from SemEval-2014 Task 4 and the “restaurants” domain (Pontiki et al., 2014) (ABSA Restaurants). We merge the first two ABSA subtasks, aspect term extraction (SB1) and aspect term polarity (SB2), into one SL task. We tokenized the original dataset with spaCy (Honnibal et al., 2020) and matched given character spans of aspect terms with token spans to create IOB2 tags for sentences (13 aspect terms could not be matched automatically in the whole dataset, so we ignored these aspect terms). Training and test sets were given. Following prior work (Wang et al., 2021), we randomly sample $10\%$ of the sentences for a validation set.
Trigger Classification.
The ACE05 dataset (Doddington et al., 2004) is a widely used trigger classification dataset. Trigger classification encapsulates two SL tasks: trigger identification, i.e., finding spans of trigger tokens and classifying those spans. We use the English train, validation, and test splits obtained with the standard ACE pre-processing tool. [^1] https://bit.ly/ace2005-preprocessing We use this tool to obtain sentences and tokens and create IOB2 tags.
Evaluation.
We evaluate all tasks with micro F1 score on IOB2 tag predictions with strict matching using “seqeval” (Nakayama, 2018), where the predicted output span must exactly match the expected output span. We evaluate only the predictions on the first token of the tokenized words from the input sequence to obtain the same number of predictions as there are target IOB2 labels.
Instruction Tuning.
A special pre-processing of each SL dataset is required to train LLMs for IT. We create instructions similarly to the ones used for NER by Wang et al. (2022b), although we require a more strict output response from decoder-based LLMs, similar to the output format used by Wang et al. (2023b). Instruction templates are created by parsing IOB2 tags to create desired outputs. Examples of instruction tuning examples created with those templates are given in Appendix A.1. To ensure fair evaluation consistent with models trained for SL, we map response spans of instruction-tuned models to IOB2 tags with various heuristics. We employ greedy span-based matching of predicted spans and their types with input tokens. If no predictions are made, all predicted spans do not align with input tokens, or if an error arises during matching due to output generation stochasticity, we treat all these scenarios as if the “O” tag was predicted for every input token. This approach is similar to the evaluation of Wang et al. (2022b).
(a) NER
(b) ABSA
(c) Trigger Classification
(d) Text Chunking
Figure 1: Micro F1 SL scores with decoder-based LLMs and different unmasking configurations (sorted by Gray code starting with all decoder layers masked – code $0$). Upper row plots show Llama2-7B model results, while lower row plots show Mistral-7B model results on validation and test sets of four SL tasks (left to right, one dataset per task). All results are averages of five runs with different seeds. The shaded area represents the standard deviation.
4.2 Training Details and Hyperparameters
To fight out-of-memory issues, we pre-process all SL datasets to have a maximum tokenized sequence length of $128$. This, however, poses no problem because there are less than $10$ sentences for each dataset and split that end up truncated independent of the tokenizer used. We use the end-of-sequence token for the models with no padding token to pad the sequences to the longest example in the batch. We randomly sample examples for training depending on the seed. For MLM-based encoder experiments, we use strong RoBERTa-base and RoBERTa-large PLMs with 125 and 355 million parameters, respectively (Liu et al., 2019). We choose “open” LLMs for unmasking experiments and IT: Llama2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), both with seven billion parameters. Since RoBERTa requires tokenization with added prefix space, we enforce the same for “open” LLMs’ tokenizers. For SL experiments, we employ a classic softmax SL head on top of the pre-trained models. Models and their implementations are taken from “Hugging Face Transformers” Wolf et al. (2020). If the SL head is not implemented, we add our implementation as it was implemented for RoBERTa models.
We use quantized LoRA (QLoRA) (Hu et al., 2021; Dettmers et al., 2023) for fine-tuning all models to ensure a fair comparison between the smallest and the largest models and make fine-tuning possible under constrained computing requirements. LoRA is applied to query and value attention matrices inside each encoder or decoder block with a fixed rank of $r=64$, a scaling parameter of $\alpha=16$, and a dropout probability of $p=0.1$. This way, only decomposed query and value matrices, and the SL head parameters are optimized with cross-entropy loss, yielding a drastic trainable parameter reduction per model. The cross-entropy loss is adjusted to consider only the first token of each tokenized word from the input sequence. The models are trained in bfloat16 precision, with loaded pre-trained weights in 4-bit NormalFloat data type, and we use double quantization. With QLoRA and the mentioned quantization configuration, all models can fit into 40GB of GPU memory of “Ampere A100” and are trained with a consistent batch size of $16$ per experiment. We train the models with paged 8-bit AdamW (Loshchilov and Hutter, 2017) optimizer to handle the memory spikes Dettmers et al. (2023). The parameters of AdamW are fixed to $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-5},\lambda=0.1$. For the learning rate scheduler, we choose cosine annealing scheduler (Loshchilov and Hutter, 2016). We apply gradient clipping to $1.0$, gradient accumulation with four steps, and gradient checkpointing. We use a consistent learning rate of $2\mathrm{e}{-4}$ across all experiments and fine-tune PLMs or LLMs over a fixed number of five epochs. All results are averages of five runs with different seeds ($seeds=[120,121,122,123,124]$), and we always pick the last model for each seed. We do not tune the hyperparameters due to the computing constraints but fix them to the mentioned recommended values in advance.
We use the PLMs and LLMs with CLM heads from “Hugging Face Transformers” for IT experiments. Training is also performed with QLoRA based on cross-entropy loss. All hyperparameters are identical to SL experiments except for the batch size. The batch size depends on the maximum sequence length for packing the dataset examples. Short examples are packed together in the same input to increase training efficiency. This length is set to $512$ for the ABSA Restaurants dataset and to $1024$ for other datasets since ABSA Restaurants has shorter instruction examples on average compared to other datasets. It also has the lowest number of examples in the training set (cf. Table 1). We observe that ABSA models need double the epochs ($10$ epochs) to learn to extract aspect terms and their polarity, and we hypothesize that this is due to the low number of training examples compared to other datasets which successfully solve the tasks with five epochs of supervised instruction fine-tuning. We generate the outputs for evaluation with default generation settings for the Llama2-7B model. We only decrease the total maximum length of the input instruction prompt with newly generated tokens to $1024$ to speed up the generation. The same generation config is used for Mistral-7B.
5 Results
Model CoNLL03 NER ABSA Restaurants ACE05 Trigger Clf. CoNLL03 Chunking Llama2-7B-SL Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Unmask Config. $0$ 0.788 0.744 0.999 0.430 0.482 0.994 0.591 0.574 0.806 0.875 0.864 0.998 $1$ 0.936 0.908 0.345 0.402 0.635 0.584 0.892 0.880 Best 0100 - - 0.488 0.546 - - - - 0101 - - - - - - - 0.885 - 0110 0.947 0.920 - - - - 0.898 - 0111 - - - - 0.653 - - - 1100 - - - - - 0.616 - - Mistral-7B-SL Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Unmask Config. $0$ 0.789 0.756 0.999 0.484 0.553 0.983 0.607 0.584 0.899 0.882 0.872 0.994 $1$ 0.953 0.927 0.581 0.641 0.668 0.626 0.904 0.892 Best 0110 - - 0.597 0.659 - - - - 0111 0.956 - - - - - - - 1101 - - - - - 0.639 - - 1110 - 0.927 - - 0.676 - - - RoBERTa-base-SL 0.897 0.883 - 0.313 0.369 - 0.609 0.508 - 0.889 0.877 - RoBERTa-large-SL 0.924 0.900 - 0.403 0.474 - 0.698 0.628 - 0.891 0.877 - Llama2-7B-IT 0.778 0.771 - 0.523 0.608 - 0.375 0.347 - 0.833 0.818 - Mistral-7B-IT 0.897 0.887 - 0.646 0.733 - 0.477 0.461 - 0.873 0.860 -
Table 2: Validation and test micro F1 SL scores for Llama2-7B and Mistral-7B models with different unmasking configurations ($0$, $1$, and other configurations which surpass the F1 score of $1$ configuration over four SL datasets – denoted as “Best”) are in the upper part of the table. The results for SL RoBERTa baselines and instruction tuning (IT) baselines are given in the lower part of the table. The best results by dataset and evaluation set are in bold. For Llama2-7B and Mistral-7B, we report the Pearson correlation coefficient between validation and test unmasking configurations ($\rho$). All results are averages of five runs with different seeds.
5.1 Layer Group Unmasking Results
Figure 1 shows the validation and test set performance of Llama2-7B and Mistral-7B for different unmasking configurations and four SL tasks. We observe some interesting phenomena. (1) In most cases, removing the CM in as little as one layer group significantly boosts the micro F1 SL score compared to keeping the CM in all layers (code $0000$) during fine-tuning. (2) It is rarely the case that removing CM from all layer groups (code $1111$), compared to removing it only from some layer groups, yields the highest score on validation and test set. (3) Depending on the task, the boosts from the CM removal can vary (highest boosts for NER, lowest for text chunking) and deviate from minimal (NER) to substantial (trigger classification). (4) Mistral-7B shows the superiority over Llama2-7B model across tasks and best configurations. (5) Performance on validation and test sets highly correlates across unmasking configurations and datasets for a fixed model.
Unmasking gains are the highest for the NER task, and the results are generally the cleanest for CoNLL03 data (NER and text chunking tasks). Refining the layer-wise CM removal inside individual layers could result with even larger micro F1 score boosts. The performance of Llama2-7B on the ABSA Restaurants dataset shows a peculiar trend where better performance is achieved when CM is kept in all layer groups compared to CM removal. It is as if allowing the model to see a broader context hurts the performance. This trend is not present in any other dataset and model combination. Furthermore, the ABSA Restaurants dataset is the only one with higher scores on the test set than on the validation set, which could be attributed to taking a random sample of training data as validation set. The deviations of results across evaluation sets are the biggest on ABSA and trigger classification tasks which could be due to the overall low number of examples per evaluation set for those tasks ($<1000$).
5.2 Comparison With Baselines
To compare the results between different unmasking configurations and strong encoder or instruction-tuned baselines, we present the F1 scores in Table 2. The results are compared against MLM-based RoBERTa PLMs fine-tuned for SL task and instruction-tuned Llama2-7B and Mistral-7B LLMs. We compare the standard $0000$ and $1111$ unmasking configurations with the best configurations overall per model and task. We also report the Pearson correlation coefficient $\rho$ between all validation and test unmasking configurations. For all SL tasks and the Llama2-7B model, we observe consistent improvement over the $1111$ unmasking configuration. For Mistral-7B, similar holds, except for the chunking tasks where $1111$ unmasking configurations yield the highest F1 scores. IT is most beneficial for the ABSA task, which aligns with findings from Scaria et al. (2023) and outperforms any unlocking configuration on validation and test sets. RoBERTa-large baseline surpassed the performance of all other models on the ACE05 validation set but failed to do so on the test set. RoBERTa baselines generally achieve high scores on SL tasks, except for ABSA. Trainin with QLoRA in combination with strategic layer unmasking creates high-performing and compact models, requiring a small number of additional parameters trained for each SL task. The ACE05 dataset exhibits the lowest correlation between performance on the validation and test sets. Nevertheless, the generally high correlation across all datasets indicates that the validation set can effectively be utilized to select the optimal unmasking configuration for each model and task.
5.3 Comparison With SOTA IE Models
Our results are competitive with the SOTA approaches. However, it is challenging to ensure a fair comparison due to differences in training and evaluation. For example, the SOTA result for ConLL03 NER is the F1 score of $94.6$, reported by Wang et al. (2021). However, this result is obtained using a model trained on the merged training and validation set. Moreover, important details of the evaluation setup were not specified (micro vs. macro F1 score, token- vs. span-based evaluation, and evaluating the model prediction on the first token of tokenized words from the input sequence vs. on all tokens). Furthermore, the authors whose work we build upon, Li et al. (2023), report SOTA results on CoNLL03 NER by removing the CM from LLama2. The score they report is $93.19$, which is competitive with Wang et al. (2021), but not SOTA. Upon code inspection, we found out that they do not use QLoRA, so they cut off sentences longer than $64$ tokens to fit the data and model into GPU memory. In contrast, we managed to keep the sequence length at $128$ tokens and cut only a few sentences per task ($<10$). These decisions affect the overall performance. Scaria et al. (2023) achieve SOTA on ABSA Restaurants dataset, reaching F1 score of $92.76$ on aspect term extraction with IT. However, they do not report the results on combined aspect term extraction and polarity but separately treat the two tasks. The SOTA F1 score results of $69.8$ on ACE05 for trigger classification are achieved with the DeepStruct model Wang et al. (2022a).
5.4 Pre-training Small Encoders and Decoders
We inspect whether removing the CM from small CLM-based decoder PLMs (under 100 million parameters) exhibits the same trend Li et al. (2023) observed for high-scale Llama2 models with multiple billion parameters. Furthermore, we verify whether MLM-based pre-training is more appropriate for reaching higher scores on the downstream SL tasks when the number of parameters, pre-training data, pre-training steps, and all other hyperparameters are fixed between encoders and decoders. We randomly initialize small LMs with four encoder or decoder blocks based on RoBERTa-base architecture. We attach a language modeling head to these small LMs and use a randomly initialized embeddings matrix based on RoBERTa-base’s vocabulary. We inherit all other parameters of the RoBERTa-base official configuration, e.g., the hidden size dimensionality of $768$. Initializing this model produces a small LM with $68$ million parameters. We use RoBERTa-base tokenizer with added prefix space to pre-tokenize the “BookCorpus” dataset (Zhu et al., 2015) consisting of $74$ million sentences. We employ this dataset to pre-train our small LMs. “BookCorpus” sentences are grouped to form blocks of size $512$ for either MLM-based pre-training of small encoder or CLM-based pre-training of small decoder LM. MLM probability is set at $0.15$. Small encoder and decoder are trained with AdamW optimizer, a cosine annealing learning rate scheduler, eight gradient accumulation steps, and bfloat16 precision with a batch size of $64$ and a learning rate of $2\mathrm{e}{-4}$. The parameters of AdamW are fixed to $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-5},\lambda=0.1$ and we apply gradient clipping to $1.0$. We save the model weights immediately after random initialization and then at five equally spaced intervals throughout each epoch. The training process continues for a pre-determined fixed number of $10$ epochs (roughly 200k steps). We train both models with same seed. Then, we load the weights of each saved checkpoint, replace the LM head with the SL head for the appropriate task, and fine-tune the models for a fixed number of five epochs with a batch size of $16$. We do this for each of the $51$ saved checkpoints with either a pre-trained encoder, a pre-trained decoder with CM kept during fine-tuning, or a pre-trained decoder with CM removed during fine-tuning. We evaluate these variants on validation sets of the four SL datasets. The results are in Figure 2. All results are averages of five fine-tuning runs with different seeds ($seeds=[120,121,122,123,124]$), and we always pick the last model for each seed. The results reveal that removing the CM on this small scale yields no gains. On average, the model without CM pefroms worse than a fine-tuned, pre-trained decoder with CM.
Moreover, small encoders struggle to keep up with decoders until around the twentieth checkpoint (fourth pre-training epoch), when they suddenly start having significantly more success on all SL tasks except trigger classification on ACE05. Training with MLM for over four epochs on the “BookCorpus” dataset drastically hurts performance on ACE05. The score RoBERTa-base achieves on ACE05 (cf. Table 2) is not much higher than the best small encoder. The gains from further MLM pre-training on ACE05 are less accentuated than the overall number of model parameters. This can be corroborated by the fact that RoBERTa-base and RoBERTa-large were both trained for 500k steps (Liu et al., 2019), and RoBERTa-large is drastically better on ACE05.
Figure 2: Validation set micro F1 SL scores of small MLM-based encoder and CLM-based decoder models (pre-trained from scratch under controlled conditions) after fine-tuning on SL tasks training data starting from specific language modeling checkpoint. “Decoder Unmask” means that the model was pre-trained with CLM objective and a CM but fine-tuned without the CM. The shaded area represents the standard deviation.
6 Conclusion
The landscape of PLMs has undergone a significant transformation, shifting from encoder to decoder dominance in various NLU tasks. This shift is attributed to the scale of decoders and their emergent zero- and few-shot capabilities. While large decoder models like ChatGPT showcase remarkable performance even without fine-tuning, challenges persist in IE-related tasks. “Open” LLMs, such as Llama2 and Mistral, introduce transparency and accessibility, facilitating experimentation with parameter-efficient fine-tuning techniques. Notably, removing the CM during fine-tuning with QLoRA enhances performance on IE tasks modeled as SL, with crucial decisions regarding layer-wise CM removal contributing to substantial improvements.
Future research should delve into refining bidirectional information flow within decoder blocks, addressing task-specific nuances, and predicting the impact of CM removal from specific layer groups without requiring extensive fine-tuning. Comparisons with encoder-based models underscore the significance of architecture and parameter scale in SL tasks. A noticeable gap exists between small-scale encoder and decoder models, where decoder models do not benefit from CM removal at a smaller scale. The ongoing upscaling of decoder models suggests a potential turning point where they might outperform encoders on SL tasks without requiring fine-tuning, leveraging their emerging properties. However, the preferred approach for now involves fine-tuning with a specialized SL head to achieve optimal performance.
Appendix A Appendix
A.1 Instruction Tuning Details
Task (dataset) Instruction example for training Instruction example for evaluation NER (CoNLL03) ### Instruction: please extract named entities and their type from the input sentence, all entity types are in options ### Options: person, location, organization, miscellaneous ### Sentence: " What we have to be extremely careful of is how other countries are going to take Germany ’s lead , " Welsh National Farmers ’ Union ( NFU ) chairman John Lloyd Jones said on BBC radio . ### Response: Germany:location;Welsh National Farmers ’ Union:organization;NFU:organization;John Lloyd Jones:person;BBC radio:organization ### Instruction: please extract named entities and their type from the input sentence, all entity types are in options ### Options: person, location, organization, miscellaneous ### Sentence: " What we have to be extremely careful of is how other countries are going to take Germany ’s lead , " Welsh National Farmers ’ Union ( NFU ) chairman John Lloyd Jones said on BBC radio . ### Response: Aspect Term Extraction and Polarity (ABSA Restaurants) ### Instruction: please extract aspect terms and their polarity from the input sentence, all polarity types are in options ### Options: positive, negative, neutral, conflict ### Sentence: The lobster sandwich is \$ 24 and although it was good it was not nearly enough to warrant that price . ### Response: lobster sandwich:conflict;price:negative ### Instruction: please extract aspect terms and their polarity from the input sentence, all polarity types are in options ### Options: positive, negative, neutral, conflict ### Sentence: The lobster sandwich is \$ 24 and although it was good it was not nearly enough to warrant that price . ### Response: Trigger Classification (ACE05) ### Instruction: please extract events and their types from the input sentence, all event types are in options ### Options: merge organization, start organization, declare bankruptcy, end organization, grant pardon, extradite, execute, impose fine, conduct trial hearing, issue sentence, file appeal, convict, file lawsuit, release on parole, arrest and send to jail, charge and indict, acquit, participate in protest or demonstration, attack, contact via written or telephone communication, meet, start position, elect, end position, nominate, transfer ownership, transfer money, marry, divorce, be born, die, sustain injury, transport ### Sentence: In his previous letter home , Apache pilot Joe Bruhl did n’t tell his family the full details about his first combat mission . ### Response: tell:contact via written or telephone communication;combat:attack ### Instruction: please extract events and their types from the input sentence, all event types are in options ### Options: merge organization, start organization, declare bankruptcy, end organization, grant pardon, extradite, execute, impose fine, conduct trial hearing, issue sentence, file appeal, convict, file lawsuit, release on parole, arrest and send to jail, charge and indict, acquit, participate in protest or demonstration, attack, contact via written or telephone communication, meet, start position, elect, end position, nominate, transfer ownership, transfer money, marry, divorce, be born, die, sustain injury, transport ### Sentence: In his previous letter home , Apache pilot Joe Bruhl did n’t tell his family the full details about his first combat mission . ### Response: Text Chunking (CoNLL03) ### Instruction: please extract chunks and their type from the input sentence, all chunk types are in options ### Options: noun phrase, verb phrase, prepositional phrase, adverb phrase, subordinated clause, adjective phrase, particles, conjunction phrase, interjection, list marker, unlike coordinated phrase ### Sentence: Rare Hendrix song draft sells for almost \$ 17,000 . ### Response: Rare Hendrix song draft:noun phrase;sells:verb phrase;for:prepositional phrase;almost \$ 17,000:noun phrase ### Instruction: please extract chunks and their type from the input sentence, all chunk types are in options ### Options: noun phrase, verb phrase, prepositional phrase, adverb phrase, subordinated clause, adjective phrase, particles, conjunction phrase, interjection, list marker, unlike coordinated phrase ### Sentence: Rare Hendrix song draft sells for almost \$ 17,000 . ### Response:
Table 3: Instruction tuning examples from the training sets of the four SL datasets.